Compute depreciation expense for all years (assume straight-line depreciation) for Bubble Corporation.

4/1/16 Purchased machine for $22,000 with a salvage value of $2,000. The estimated useful life is 20 years.

7/1/17 $2,000 improvement to machine. Salvage value changed to $250

7/1/18 Changed estimated useful life to a total of 10 years remaining

11/1/19 Sold machine for $20,000

Please do the years 2016, 2017, 2018, and 2019.

The answers are

$750

$1,100

$2,190

$1,825

I just am confused how to solve the problem. Please help
Answer #1
Straight-line rate
Years Annual Depreciation Expense
2016 $750
2017 $1,100
2018 $1,665
2019 $1,775
Depreciation per year = ($22,000 - $2,000)/240 months = $83.33
For 2016 = $83.33 x 9 months = $750
Cost $22,000
Less: Depreciation for 2016 $750
$21,250
Less: Depreciation for 2017 - 6 months (till 6/30/17)
($83.33 x 6 months) $500
$20,750
Add: Improvement $2,000
$22,750
Less: Depreciation for 2017 - 6 months
($22,750 - $250)/225 months = $100 per month $600
$22,150
Less: Depreciation for 2018 - 6 months
($100 x 6 months) $600
$21,550
Less: Depreciation for 2018 - 6 months
($21,550 - $250)/120 months = $177.5 per month $1,065
$20,485
Less: Depreciation for 2019 - 10 months
($177.5 x 10 months) $1,775
$18,710